# Almost equivalences between Tamarkin category and Novikov sheaves

### v2

@@ 3.1 Derived completeness
    $\bigoplus_{\mathbb{N}}\Lambda_{0}$ is not complete. The completion is denoted by ^$\widehat{\bigoplus}_{\mathbb{N}}\Lambda_{0}$. Concretely, it consists of a sequence $(x_{i})_{i\in\mathbb{N}}$ of $\Lambda_{0}$ satisfying $\lim_{i\rightarrow\infty}v(x_{i})=\infty$ for the valuation $v$.
@@
    (^$(\widehat{\bigoplus}_{\mathbb{N}}\Lambda_{0})\otimes_{\Lambda_{0}}(\widehat{\bigoplus}_{\mathbb{N}}\Lambda_{0})$ is not complete. Let us write the basis explicitly: (^$(\widehat{\bigoplus}_{i\in\mathbb{N}}\Lambda_{0}e_{i})\otimes_{\Lambda_{0}}(\widehat{\bigoplus}_{j\in\mathbb{N}}\Lambda_{0}f_{j})$. For example, in the completion, we have $\sum_{i=0}^{\infty}T^{i}e_{i}\otimes f_{i}$, but is not in (^$(\widehat{\bigoplus}_{\mathbb{N}}\Lambda_{0})\otimes_{\Lambda_{0}}(\widehat{\bigoplus}_{\mathbb{N}}\Lambda_{0})$.
@@
$\varphi\colon\widehat{\bigoplus}_{n\in\mathbb{N}}\Lambda_{0}\rightarrow\prod_{n\in\mathbb{N}}\Lambda_{0},(x_{1},x_{2},x_{3},...)\mapsto(x_{1},Tx_{2},T^{2}x_{3},...).$ (3.2)
in the full subcategory of the complete modules in $\mathrm{Mod}^{\heartsuit}(\Lambda_{0}).$ We would like to check the homomorphism theorem. We first compute the coimage. The coimage is defined by the cokernel of the map $\ker(\varphi)\rightarrow\widehat{\bigoplus}_{n\in\mathbb{N}}\Lambda_{0}$. Hence it is isomorphic to ^$\widehat{\bigoplus}_{n\in\mathbb{N}}\Lambda_{0}$, since $\varphi$ is injective.
@@
Let $A$ be a ring and $I$ is a finitely generated ideal of $A$. We say an object $M\in\mathrm{Mod}(A)$ is derived complete if $\operatorname{Hom}_{\mathrm{Mod}(A)}(A[f^{-1}],M)\cong 0$ for any $f\in I$.
@@
We denote the subcategory of derived complete modules of $\mathrm{Mod}(\Lambda_{0})$ by $\mathrm{Mod}_{\mathrm{c}}(\Lambda_{0})$. We also set
@@
    $\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})$ is a presentable category. We denote the coproduct (resp. monoidal operation) in $\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})$ by $\widehat{\bigoplus}$ (resp. $\widehat{\otimes}$).
@@
The inclusion is obviously product-preserving, hence we have a left adjoint. The coproduct is given by
^$\widehat{\bigoplus}_{i\in I}\mathcal{E}_{i}:=\widehat{\bigoplus_{i\in I}\mathcal{E}_{i}}.$ (3.5)
@@
    ^$\widehat{\bigoplus}_{n\in\mathbb{N}}\Lambda_{0}\rightarrow\prod_{n\in\mathbb{N}}\Lambda_{0},(x_{1},x_{2},x_{3},...)\mapsto(x_{1},Tx_{2},T^{2}x_{3},...)$ (3.6)
@@ 4 Almost modules
4.1 Almost isomorphism
We have the usual Novikov ring $\Lambda_{0}$ and its maximal ideal $\mathfrak{m}$.
@@ 4.1 Almost isomorphism
There exists a right adjoint [GR03] given by
@@
4.2 Almost isomorphisms
Let $\mathcal{C}$ be a $\Lambda_{0}$-linear category. We denote the $\Lambda_{0}$-linear Yoneda embedding by $\mathcal{Y}\colon\mathcal{C}\rightarrow\mathrm{Mod}(\mathcal{C},\Lambda_{0})$.
@@ 4.2 Almost isomorphisms
    Let $\mathrm{Mod}(\mathcal{C},\Lambda_{0})$ be the category of $\Lambda_{0}$-modules. An almost zero module $M$ is a module such that $\mathfrak{m}\otimes M\cong 0$. In other words, $M(c)$ is almost zero for any $c\in\mathcal{C}$.
@@
    We say $\mathcal{E},\mathcal{F}\in\mathcal{C}$ are almost isomorphic to $\mathcal{Y}(\mathcal{E})$ and $\mathcal{Y}(\mathcal{F})$ are almost isomorphic. In this case, we denote it by $\mathcal{E}\cong_{\mathrm{a}}\mathcal{F}$.
@@
By Gabber–Ramero, the module ~$\widetilde{}\mathfrak{m}:=\Lambda_{+}\otimes_{\Lambda_{0}}\Lambda_{+}$ is flat. We then have an almost isomorphism
~$\widetilde{}\mathfrak{m}\otimes_{\Lambda_{0}}\mathcal{Y}(\operatorname{Cone}(f^{\prime}))\rightarrow\mathcal{Y}(\operatorname{Cone}(f)).$ (4.6)
We have
$\mathcal{Y}(\operatorname{Cone}(f^{\prime}))\cong_{\mathrm{a}}\widetilde{}\mathfrak{m}\otimes_{\Lambda_{0}}\mathcal{Y}(\operatorname{Cone}(f^{\prime}))\cong_{\mathrm{a}}\mathcal{Y}(\operatorname{Cone}(f)).$ (4.7)
@@
$\mathcal{Y}\colon\mathcal{C}\rightarrow\mathrm{Mod}(\mathcal{C},L_{0}^{\mathbb{G}}):=\mathop{\mathrm{\bf{Fun}}}(\mathcal{C}^{op},\mathrm{Mod}(L_{0}^{\mathbb{G}}));\mathcal{E}\mapsto\operatorname{Hom}(-,\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}\mathcal{E}).$ (4.9)
@@
    We say $\mathcal{E},\mathcal{F}\in\mathcal{C}$ are almost isomorphic to $\mathcal{Y}(\mathcal{E})$ and $\mathcal{Y}(\mathcal{F})$ are almost isomorphic. In this case, we denote it by $\mathcal{E}\cong_{\mathrm{a}}\mathcal{F}$.
@@ 4.3 Almost equivalence
Let $\mathcal{C}_{1},\mathcal{C}_{2}$ be categories over $L_{0}^{\mathbb{G}}$.
@@
    We say $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are almost equivalent (, or $\mathcal{C}_{1}\cong_{\mathrm{a}}\mathcal{C}_{2}$) if there exists a morphism from $\mathcal{C}_{1}$ to $\mathcal{C}_{2}$ such that $f$ is almost fully faithful and almost essentially surjective.
@@
We consider the category $\mathrm{Mod}(L_{0}^{\mathbb{G}})$. Since an object of $\mathrm{Mod}(L_{0}^{\mathbb{G}})$ carries a $\mathbb{R}/\mathbb{G}$-grading, we can shift it. We denote the resulting shift functor by $T_{c}$. Then, for any $M,N\in\mathrm{Mod}(L_{0}^{\mathbb{G}})$, we have
@@
which is an $L_{0}^{\mathbb{G}}$-module.
@@
The category $\mu^{\mathbb{G}}(M)$ will be introduced in the next section. We have shift operations $T_{c}$ parmetrized by $c\in\mathbb{R}/\mathbb{G}$. Then, for any $\mathcal{E},\mathcal{F}\in\mu^{\mathbb{G}}(M)$, we have
$\operatorname{Hom}_{\mu^{\mathbb{G}}(M)}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}\mathcal{E},\mathcal{F}),$ (4.13)
which is an $L_{0}^{\mathbb{G}}$-module.
@@ 5.1 Basics
We consider the addition action of a subgroup $\mathbb{G}\subset\mathbb{R}$ on $\mathbb{R}_{t}$ as a discrete group action. Then we can consider the derived category of equivariant $\mathbb{K}$-module sheaves $\mathrm{Sh}^{\mathbb{G}}(\mathbb{R}_{t},\mathbb{K})$.
We denote the subcategory spanned by the object whose microsupport contained in ×$\mathbb{R}\times\mathbb{R}_{\leq 0}\subset\mathbb{R}\times\mathbb{R}\cong T^{*}\mathbb{R}_{t}$ by $\mathrm{Sh}^{\mathbb{G}}_{\mathbb{R}_{\leq 0}}(\mathbb{R}_{t},\mathbb{K})$. We set
@@
    $\mu^{\mathbb{G}}(*)$ has a monoidal structure defined by the convolution product.
@@
We have an almost isomorphism of almost $L_{0}^{\mathbb{G}}$-modules
@@
We first consider the following exact triangle:
@@
Since the $i$-th cohomologies vanish for $i>1$, we will only take care of $H^{1}\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})$. (Note that $H^{1}\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\cong H^{1}\operatorname{Hom}(\mathbb{K}_{t<0},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})$.)
@@
$\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{\mathbb{R}})\simeq\bigoplus_{c\in\mathbb{R}}\mathbb{K}.$ (5.5)
Hence we have $H^{1}\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\simeq 0$. Hence
@@
for any $a>0$. This implies that the action of $T^{a}$ on (5.6) is zero for any $a>0$. Hence it is almost zero. This completes the proof. ∎
@@
$\operatorname{End}_{\mu^{\mathbb{G}}(*)}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu})\cong L_{0}^{\mathbb{G}}.$ (5.8)
@@
$\begin{split}\operatorname{End}_{\mu^{\mathbb{G}}(*)}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu})\cong\prod_{c\in\mathbb{R}/\mathbb{G}}\operatorname{Hom}_{\mu^{\mathbb{O}}(*)}(\mathbb{K}_{t\geq 0},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\cong\prod_{c\in\mathbb{R}/\mathbb{G}}\Lambda_{0}\cong L_{0}^{\mathbb{G}}\end{split}$ (5.9)

### v3

@@ 3.1 Derived completeness
    $\bigoplus_{\mathbb{N}}\Lambda_{0}$ is not complete. The completion is denoted by $\widehat{\bigoplus}_{\mathbb{N}}\Lambda_{0}$. Concretely, it consists of a sequence $(x_{i})_{i\in\mathbb{N}}$ of $\Lambda_{0}$ satisfying $\lim_{i\rightarrow\infty}v(x_{i})=\infty$ for the valuation $v$.
@@
    $(\widehat{\bigoplus}_{\mathbb{N}}\Lambda_{0})\otimes_{\Lambda_{0}}(\widehat{\bigoplus}_{\mathbb{N}}\Lambda_{0})$ is not complete. Let us write the basis explicitly: $(\widehat{\bigoplus}_{i\in\mathbb{N}}\Lambda_{0}e_{i})\otimes_{\Lambda_{0}}(\widehat{\bigoplus}_{j\in\mathbb{N}}\Lambda_{0}f_{j})$. For example, in the completion, we have $\sum_{i=0}^{\infty}T^{i}e_{i}\otimes f_{i}$, but that is not in $(\widehat{\bigoplus}_{\mathbb{N}}\Lambda_{0})\otimes_{\Lambda_{0}}(\widehat{\bigoplus}_{\mathbb{N}}\Lambda_{0})$.
@@
[…]
in the full subcategory of the complete modules in $\mathrm{Mod}^{\heartsuit}(\Lambda_{0}).$ We would like to check the homomorphism theorem. We first compute the coimage. The coimage is defined by the cokernel of the map $\ker(\varphi)\rightarrow\widehat{\bigoplus}_{n\in\mathbb{N}}\Lambda_{0}$. Hence it is isomorphic to $\widehat{\bigoplus}_{n\in\mathbb{N}}\Lambda_{0}$, since $\varphi$ is injective.
@@
Let $A$ be a ring and $I$ is a finitely generated ideal of $A$. We say an object $M\in\mathrm{Mod}(A)$ is derived complete if $\operatorname{Hom}_{\mathrm{Mod}(A)}^{\bullet}(A[f^{-1}],M)\cong 0$ for any $f\in I$.
@@
We denote the subcategory of derived complete modules with respect to $I=(T^{1})$ of $\mathrm{Mod}(\Lambda_{0})$ by $\mathrm{Mod}_{\mathrm{c}}(\Lambda_{0})$. We also set
@@
    $\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})$ is a stable and presentable category. We denote the coproduct in $\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})$ by $\widehat{\bigoplus}$.
@@
Note that the derived complete objects is characterized by the right orthogonal to the homotopy limit of the sequence
$N_{c}:=\cdots\xrightarrow{T^{c}}\Lambda_{0}\xrightarrow{T^{c}}\Lambda_{0}\xrightarrow{T^{c}}\Lambda_{0}.$ (3.5)
for some $c>0$ (and any $c>0$). We also set
$N_{c}^{\mathbb{G}}:=\cdots\xrightarrow{T^{c}}L^{\mathbb{G}}_{0}\xrightarrow{T^{c}}L^{\mathbb{G}}_{0}\xrightarrow{T^{c}}L^{\mathbb{G}}_{0}.$ (3.6)
for $c\in\mathbb{G}\cap\mathbb{R}_{>0}$. It is easy to see that the vanishing of $\operatorname{Hom}_{\mathrm{Mod}(L_{0}^{\mathbb{G}})}(N_{c}^{\mathbb{G}},M)$ is equivalent to the vanishing of $\operatorname{Hom}_{\mathrm{Mod}(\Lambda_{0})}(N_{c},\mathfrak{f}(M))$. Hence $\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})$ is the full subcategory of the $S$-local objects where $S=\left\{N_{c}^{\mathbb{G}}\right\}_{c\in\mathbb{G}\cap\mathbb{R}_{>0}}$. Since $\mathrm{Mod}(L_{0}^{\mathbb{G}})$ is compactly generated, [Lur09, 5.5.4.15] implies that $\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})$ is presentable and the inclusion $\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})\hookrightarrow\mathrm{Mod}(L_{0}^{\mathbb{G}})$ admits a left adjoint. Now it is easy to see the left adjoint given by the formula (3.4), and the coproduct is given by
$\widehat{\bigoplus}_{i\in I}\mathcal{E}_{i}:=\widehat{\bigoplus_{i\in I}\mathcal{E}_{i}}$. (3.7)
@@
[…]
@@ 4 Almost modules
4.1 Almost isomorphism in $\mathrm{Mod}(R)$
We have the usual Novikov ring $\Lambda_{0}$ and its maximal ideal $\mathfrak{m}$. Note that $\mathfrak{m}$ is flat.
@@ 4.1 Almost isomorphism in $\mathrm{Mod}(R)$
There exists a right adjoint [GR03] denoted by
@@
4.2 Almost isomorphisms between $\mathcal{C}$-modules
Let $\mathcal{C}$ be a stable $\Lambda_{0}$-linear category.
@@ 4.2 Almost isomorphisms between $\mathcal{C}$-modules
    Let $\mathrm{Mod}(\mathcal{C},\Lambda_{0})$ be the category of $\mathcal{C}$-modules over $\Lambda_{0}$. An almost zero module $M$ is a module such that $\mathfrak{m}\otimes M\cong 0$. In other words, $M(c)$ is almost zero for any $c\in\mathcal{C}$.
@@
    Two objects $M,N$ in $\mathrm{Mod}(\mathcal{C},\Lambda_{0})$ are said to be almost isomorphic if they are isomorphic in $\mathrm{Mod}(\mathcal{C},\Lambda_{0}^{\mathrm{a}})$.
  4. 4.
    We denote the $\Lambda_{0}$-linear Yoneda embedding by $\mathcal{Y}\colon\mathcal{C}\rightarrow\mathrm{Mod}(\mathcal{C},\Lambda_{0})$. We say $\mathcal{E},\mathcal{F}\in\mathcal{C}$ are almost isomorphic if $\mathcal{Y}(\mathcal{E})$ and $\mathcal{Y}(\mathcal{F})$ are almost isomorphic. In this case, we denote it by $\mathcal{E}\cong_{\mathrm{a}}\mathcal{F}$.
@@
Recall that $\mathfrak{m}$ is flat. Then
$\mathfrak{m}\otimes_{\Lambda_{0}}\mathcal{Y}(\operatorname{Cone}(f))=\operatorname{Cone}(\mathfrak{m}\otimes_{\Lambda_{0}}\mathcal{Y}(f)\colon\mathfrak{m}\otimes_{\Lambda_{0}}\mathcal{Y}(\mathcal{E})\to\mathfrak{m}\otimes_{\Lambda_{0}}\mathcal{Y}(\mathcal{F})).$ (4.6)
Since $f-f^{\prime}$ is almost zero, $\mathfrak{m}\otimes_{\Lambda_{0}}\mathcal{Y}(f)=\mathfrak{m}\otimes_{\Lambda_{0}}\mathcal{Y}(f^{\prime})$. Hence we have an isomorphism
$\mathfrak{m}\otimes_{\Lambda_{0}}\mathcal{Y}(\operatorname{Cone}(f^{\prime}))\cong\mathfrak{m}\otimes_{\Lambda_{0}}\mathcal{Y}(\operatorname{Cone}(f)).$ (4.7)
Since $\mathfrak{m}$ is almost isomorphic to $\Lambda_{0}$, we have
$\mathcal{Y}(\operatorname{Cone}(f^{\prime}))\cong_{\mathrm{a}}\mathfrak{m}\otimes_{\Lambda_{0}}\mathcal{Y}(\operatorname{Cone}(f^{\prime}))\cong\mathfrak{m}\otimes_{\Lambda_{0}}\mathcal{Y}(\operatorname{Cone}(f))\cong_{\mathrm{a}}\mathcal{Y}(\operatorname{Cone}(f)).$ (4.8)
where the middle equality uses (4.7). ∎
@@
$\mathcal{Y}\colon\mathcal{C}\rightarrow\mathrm{Mod}(\mathcal{C},L_{0}^{\mathbb{G}}):=\mathop{\mathrm{\bf{Fun}}}(\mathcal{C},\mathrm{Mod}(L_{0}^{\mathbb{G}}));\mathcal{E}\mapsto\operatorname{Hom}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}\mathcal{E},-)$ (4.10)
where $\mathop{\mathrm{\bf{Fun}}}$ is the functor category.
@@
    Two objects $M,N$ in $\mathrm{Mod}(\mathcal{C},{L_{0}^{\mathbb{G}}})$ are said to be almost isomorphic if they are isomorphic in $\mathrm{Mod}(\mathcal{C},{L_{0}^{\mathbb{G}}}^{\mathrm{a}})$.
  4. 4.
    We say $\mathcal{E},\mathcal{F}\in\mathcal{C}$ are almost isomorphic if $\mathcal{Y}(\mathcal{E})$ and $\mathcal{Y}(\mathcal{F})$ are almost isomorphic. In this case, we denote it by $\mathcal{E}\cong_{\mathrm{a}}\mathcal{F}$.
@@ 4.3 Almost equivalence
Let $\mathcal{C}_{1},\mathcal{C}_{2}$ be cocomplete categories over $L_{0}^{\mathbb{G}}$.
@@
    We say $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are almost equivalent (, or $\mathcal{C}_{1}\cong_{\mathrm{a}}\mathcal{C}_{2}$) if there exists a functor $f$ from $\mathcal{C}_{1}$ to $\mathcal{C}_{2}$ such that $f$ is almost fully faithful and almost essentially surjective.
@@
We consider the category $\mathrm{Mod}(L_{0}^{\mathbb{G}})$. Since an object of $\mathrm{Mod}(L_{0}^{\mathbb{G}})$ carries an $\mathbb{R}/\mathbb{G}$-grading, we can shift it. We denote the resulting shift functor by $T_{c}$. Then, for any $M,N\in\mathrm{Mod}(L_{0}^{\mathbb{G}})$, we have
@@
which is a right $L_{0}^{\mathbb{G}}$-module.
@@
The category $\mu^{\mathbb{G}}(T^{*}X)$ will be introduced in the next section. We have shift operations $T_{c}$ parametrized by $c\in\mathbb{R}/\mathbb{G}$. Then, for any $\mathcal{E},\mathcal{F}\in\mu^{\mathbb{G}}(T^{*}X)$, we have
$\operatorname{Hom}_{\mu^{\mathbb{G}}(T^{*}X)}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}\mathcal{E},\mathcal{F}),$ (4.14)
which is a right $L_{0}^{\mathbb{G}}$-module.
@@ 5.1 Basics
We consider the addition action of a subgroup $\mathbb{G}\subset\mathbb{R}$ on $\mathbb{R}_{t}$ as a discrete group action. Then we can consider the derived $\infty$-category of equivariant $\mathbb{K}$-module sheaves $\mathrm{Sh}^{\mathbb{G}}(\mathbb{R}_{t},\mathbb{K})$. The notion of microsupport $S\SS$ is defined for each object of $\mathrm{Sh}(\mathbb{R}_{t},\mathbb{K})$. For each object $\mathrm{Sh}^{\mathbb{G}}(\mathbb{R}_{t},\mathbb{K})$, we define its microsupport by that of the image under the forgetful functor $\mathrm{Sh}^{\mathbb{G}}(\mathbb{R}_{t},\mathbb{K})\to\mathrm{Sh}(\mathbb{R}_{t},\mathbb{K})$ forgetting the equivariancy.
We denote the subcategory spanned by the object whose microsupport contained in $\mathbb{R}\times\mathbb{R}_{\leq 0}\subset\mathbb{R}\times\mathbb{R}\cong T^{*}\mathbb{R}_{t}$ by $\mathrm{Sh}^{\mathbb{G}}_{\mathbb{R}_{\leq 0}}(\mathbb{R}_{t},\mathbb{K})$. We set
@@
By Tamarkin’s cutoff functor (see [Kuw24]), the category $\mu^{\mathbb{G}}(*)$ has a canonical fully faithful embedding $\mu^{\mathbb{G}}(*)\hookrightarrow\mathrm{Sh}^{\mathbb{G}}(\mathbb{R},\mathbb{K})$. Most of the hom-computations below are done through this embedding.
@@
    $\mu^{\mathbb{G}}(*)$ has a monoidal structure $\star$ defined by the convolution product.
@@
  4. 4.
    The forgetful functor $\mu^{\mathbb{G}}(*)\rightarrow\mu^{\mathbb{O}}(*)$ forgetting the equivariancy admits left adjoint given by $\mathcal{E}\mapsto\bigoplus_{c\in\mathbb{G}}\mathbb{K}_{t\geq c}\star\mathcal{E}$ with the obvious equivariant structure on the image.
@@
We have an almost isomorphism of almost $\Lambda_{0}^{\mathbb{G}}$-modules
@@
In the rest of this proof, $\operatorname{Hom}$ means dg-Hom space in $\mathrm{Sh}(\mathbb{R},\mathbb{K})$. We first consider the following exact triangle:
@@
Since the $i$-th cohomologies vanish for $i>1$ due to [KS94, Proposition 3.2.2], we will only take care of $H^{1}\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})$.
@@
$\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{\mathbb{R}})\simeq\bigoplus_{c\in\mathbb{R}}\mathbb{K}$ (5.5)
since $\mathbb{K}_{\mathbb{R}}$ is compact in the full subcategory of constant sheaves on $\mathbb{R}$. Hence we have $H^{1}\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\simeq 0$. This together with (5.3) implies that
@@
for any $a>0$ by the non-characteristic deformation lemma. This implies that the action of $T^{a}$ on (5.6) is zero for any $a>0$. Hence it is almost zero. This completes the proof. ∎
@@
$\operatorname{End}_{\mu^{\mathbb{G}}(*)}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu})\cong_{\mathrm{a}}L_{0}^{\mathbb{G}}.$ (5.8)
@@
$\begin{split}\operatorname{End}_{\mu^{\mathbb{G}}(*)}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu})\cong_{\mathrm{a}}\prod_{c\in\mathbb{R}/\mathbb{G}}\operatorname{Hom}_{\mu^{\mathbb{O}}(*)}(\mathbb{K}_{t\geq 0},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\cong_{\mathrm{a}}\prod_{c\in\mathbb{R}/\mathbb{G}}\Lambda_{0}^{\mathbb{G}}\cong_{\mathrm{a}}L_{0}^{\mathbb{G}}\end{split}$ (5.9)